You are dealt two cards successively (without replacement) from a shuffled deck of 52 playing cards. Find the probability that the first card is a king and the second card is a queen.
Answer by sudhanshu_kmr(1152)


there are 4 king and 4 queen.
probability that first card is a king = 4/52
probability that second card is a queen = 4/51
(no replacement so only 51 will remain)

required probability = 4/52 * 4/51
= 1/13 * 4/51
= 4/663